OpenStudy (anonymous):

How to find x-intercepts of f(x) = 2x^3-x-2? I'm not really sure how to approach this question... with it being a cubic function?

ganeshie8 (ganeshie8):

you want to solve it algebraically ?

OpenStudy (anonymous):

Yeap! @ganeshie8

ganeshie8 (ganeshie8):

lets start by dividing 2 through out so that leading coeffecient becomes 1

ganeshie8 (ganeshie8):

\(2x^3-x-2 = 0\) \(x^3-x/2-1 = 0\)

hero (hero):

I might have an approach to this myself as well.

ganeshie8 (ganeshie8):

ive just made up my method but not quite too sure of it yet... please show your method @Hero

hero (hero):

I made mine up too bro, but I'll post something.

hero (hero):

\[2x^3 - 2 - x = 0\]\[2(x^3 - 1) - x = 0\]\[2((x - 1)(x^2 + x + 1)) - x = 0\]\[2(x -1)(x^2 + x + 1) - x = 0\]\[(x-1)\left(2(x^2 + x + 1)) - \frac{x}{(x-1)}\right) =0\]

hero (hero):

Now use zero product property

hero (hero):

@cmolina19 @ganeshie8

hero (hero):

There's probably a mistake in there somewhere

OpenStudy (anonymous):

Wait how did you get the second set of parentheses? o.O I've been trying to do this on my own.. and I'm actually wondering.. is it possible to do this: 2x^3-x-2=0 2x^3-x=2 x(x^2-1)=0 x=0, x=1 ???

OpenStudy (anonymous):

wait oh crap, changed the 2 to zero

hero (hero):

Your second step to third step is flawed

OpenStudy (anonymous):

ok, x(x^2-1)=2 x=2, x=sqrt(3)

OpenStudy (anonymous):

Yeah, that's right. I noticed it once I posted it lol

OpenStudy (anonymous):

Nevermind, plugging it into the calculator it's not correct

hero (hero):

I used difference of cubes to get the expression above

hero (hero):

\[x^3 - 1 = x^3 - 1^3\]

hero (hero):

But something is flawed with it since x = 1 is not a solution

hero (hero):

Do you get what I did (or tried to do)?

OpenStudy (anonymous):

I do, but not the last line :$ (x−1)(2(x2+x+1))−x(x−1))=0

hero (hero):

That's not what the last line reads.

hero (hero):

That last expression is -x/(x-1)

hero (hero):

It's what happens when you factor out x - 1

ganeshie8 (ganeshie8):

that gives back cubic again

hero (hero):

I know. You must have worked on it @ganshie8

ganeshie8 (ganeshie8):

yeah trying ur method as it looks simple but ending up wid cubic no matter wat i do

hero (hero):

x = 1 is not a solution so immediately we know something is wrong.

OpenStudy (anonymous):

Yeah, whoops. Copied and pasted it to here, so it missed some characters.

ganeshie8 (ganeshie8):

yeah that also there

ganeshie8 (ganeshie8):

try this :- 1) find one root \(k\) by doing some algebra tricks (I am going to help you with this) 2) then divide \((x-k)\) turning the cubic to quadratic 3) quadratic can be easily handled

hero (hero):

Yes, of course. That's the traditional way of doing it. I was trying to use some factoring method.

ganeshie8 (ganeshie8):

traditional is painful and lengthy also, anyways i feel like giving it a try now :)

ganeshie8 (ganeshie8):

\(\large 2x^3-x-2 = 0\) \(\large x^3-x/2-1 = 0\) --------------(1) say, x = u+v \(\large x^3 = u^3+v^3 + 3uv(u+v) \) \(\large x^3 = u^3+v^3 + 3uvx \) \(\large x^3 -3uvx - (u^3+v^3) = 0 \) -------(2)

ganeshie8 (ganeshie8):

compare x coefficient and constant terms between (1) and (2)

ganeshie8 (ganeshie8):

\(3uv = 1/2 => uv = 1/6 => u^3v^3 = 1/216\) \(u^3+v^3 = 1\)

ganeshie8 (ganeshie8):

quadratic with \(u^3, v^3\) as roots :- \(t^2-t+1/216 = 0\) here next steps :- 1) solve t by completing the square 2) taking cube root of solutions above, gives u value of u and v 3) you have x, since x = u+v

ganeshie8 (ganeshie8):

give it a try

hero (hero):

Good luck with that bro.

OpenStudy (anonymous):

wait, what is t again?

ganeshie8 (ganeshie8):

t is just an intermediate variable, solve the quadratic in t. solutions will be u^3 and v^3

ganeshie8 (ganeshie8):

\(t^2-t+1/216 = 0\) \(216(t^2-t) + 1 = 0\) \(216((t-1/2)^2) = 53\) \((t-1/2)^2 = 53/216\) solve t keep going